OpenStudy (bobdole):

x-intercept 2, y-intercept 3/5 the slope-intercept form for the line is ??? plz

OpenStudy (lgbasallote):

are you familiar with the intercept form\[\frac xa + \frac yb = 1\] where: a = x-intercept b = y-intercept

OpenStudy (bobdole):

sort of

OpenStudy (lgbasallote):

good. so try substituting. what do you get?

OpenStudy (bobdole):

would it be like y=m2+3/5?

OpenStudy (lgbasallote):

m2?

OpenStudy (bobdole):

y =m(x)+b is formula??

OpenStudy (bobdole):

this is entire question

OpenStudy (bobdole):

find the slope-intercept form for the line satisfying the following conditions. x-intercept 2, y-intercept 3/5 the slope-intercept form for the line is ??? plz

OpenStudy (lgbasallote):

it will look like that yes

OpenStudy (lgbasallote):

y = mx + b

OpenStudy (bobdole):

where would we put the 3/5 would it be x or y ?

OpenStudy (lgbasallote):

b

OpenStudy (bobdole):

y=2+3/5?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

like i said follow what i say first. you'll get there

OpenStudy (bobdole):

y=2x=3/5

OpenStudy (lgbasallote):

substitute 2 and 3/5 into \[\frac xa + \frac yb = 1\]

OpenStudy (bobdole):

2/1 + 3/5

OpenStudy (lgbasallote):

what?

OpenStudy (bobdole):

is that how i would set up the problem?

OpenStudy (bobdole):

not really grasping how to start it

OpenStudy (lgbasallote):

follow what i say....

OpenStudy (bobdole):

k

OpenStudy (lgbasallote):

substitute 2 and 3/5 into \[\frac xa + \frac yb = 1\]

OpenStudy (bobdole):

where do we get a and b?

OpenStudy (lgbasallote):

a = x-intercept b = y-intercept

OpenStudy (bobdole):

but y would be 3/5 and what would be the bottom number for 2 also?

OpenStudy (lgbasallote):

huh? a is the x-intercept and b is the y-intercept. so what are the x and y-intercepts here?

OpenStudy (bobdole):

x= 2 and y would be 3/5 so would it be 2/1 +3/5

OpenStudy (lgbasallote):

x is not 2 and y is not 3/5

OpenStudy (lgbasallote):

substitute 2 into a and substitute 3/5 into b

OpenStudy (bobdole):

x/2 + y over 3/5

OpenStudy (lgbasallote):

where did = 1 go?

OpenStudy (bobdole):

so 2/1x and 3/5 y

OpenStudy (lgbasallote):

no... x/2 + y/3/5 = 1

OpenStudy (bobdole):

i am so lost so how would the form look like or is that the form

OpenStudy (lgbasallote):

just follow what im saying and stop thinking about the answer

OpenStudy (lgbasallote):

do you get what i did?

OpenStudy (bobdole):

ok

OpenStudy (bobdole):

yes

OpenStudy (lgbasallote):

now multiply 3/5 to all the terms...what do you get?

OpenStudy (bobdole):

1 for 2/1times 3/5 and 9/5 for 3/5 times 3/5

OpenStudy (lgbasallote):

....what are you doing? you multiply 3/5 to \[\huge \frac x2 + \frac y{\frac 35} = 1\]

OpenStudy (anonymous):

let me help x-intercept is 2 so we have the point (2,0) y- intercept is 3/5 so we have (0,3/5) y=mx+b b is y intercept, so b = 3/5 y=mx+3/5 now sub (2,0) 0=m(2)+3/5 2m=-3/5 m= (-3/5)*(1/2) m=-3/10 so what is your equations

OpenStudy (bobdole):

y=3/10x +3/5

OpenStudy (anonymous):

im sure bobdole wasnt taught \[\frac{ x }{ a }+\frac{ y }{ b } method\]

OpenStudy (bobdole):

not even a little

OpenStudy (anonymous):

you should have -3/10

OpenStudy (anonymous):

another way is to look at graph do you have graph paper?

OpenStudy (bobdole):

y=-3/10x-3/5

OpenStudy (bobdole):

do not have graph paper

OpenStudy (anonymous):

no y=-3/10 x +3/5

OpenStudy (bobdole):

oh ok thanx you so very very much

OpenStudy (bobdole):

dont the bottom have to be common denominators?

OpenStudy (anonymous):

no because you are wanting the slope intercept form :-)